2.Two different single nucleotide polymorphisms in the ataxia telangiectasia mutated (ATM) gene (rs11212617) are associated with the response to the glucose-lowering drug metformin which is used in treating type 2 diabetes. These alleles vary in frequency among ethnic populations. In a particular South Indian population, the following genotype frequencies were observed. Calculate the frequencies of the A, C1 and C2 alleles in this population. '

Genotype AA AC1 C1C1 AC2 C2C2 C1C2

Frequency 0.49 0.406 0.0841 0.014 0.0001 0.0058

3. Cystic fibrosis is a recessive condition with a simple genetic basis. Cystic fibrosis sufferers result when an individual inherits two copies of a defective CFTR gene (the CFTR gene produces a protein involved in the production of sweat, digestive fluids and mucus). The data in Table 1 are from a sample of 2,987 Canadians and the functional and defective alleles are denoted CFTR and CFTR- respectively:

Genotype: CFTR/CFTR CFTR/CFTR- CFTR-/CFTR-

# of individuals: 2886 100 1

What are the frequencies of the CFTR and CFTR- alleles in this sample?
Answer #1

2.

Genotype Frequency
AA 0.49
AC1 0.406
C1C1 0.0841
AC2 0.014
C2C2 0.0001
C1C2 0.0058

The frequencies of the alleles A,C1 and C2 are

Frequency of allele A = 0.49+0.406*0.5+0.014*0.5 = 0.7

Frequency of allele C1 = 0.406*0.5+0.0841+0.0058*0.5 = 0.29

Frequency of allele C2 = 0.014*0.5+0.0001+0.058*0.5 = 0.01

alleles Frequency
A 0.7
C1 0.29
C2 0.01

3.

Genotype

number of

individuals

Frequency
CFTR/CFTR 2886 0.9662
CFTR/CFTR- 100 0.03347
CFTR-/CFTR- 1 0.00033
Total 2987 1

The frequency of the alleles CFTR and CFTR- are:

frequency of the allele CFTR = 0.9662+ 0.03347*0.5 = 0.9829

frequency of the allele CFTR- = 0.03347+0.00033*0.5 = 0.01717

Allele Frequency
CFTR 0.9829
CFTR- 0.0171